If 3x = square root of 2, what is the value of 12x^6 - 9x^4 + 6x^2?
Answer by rothauserc(4718)

3x = square root(2)
:
square both sides of =
:
9x^2 = 2
:
x^2 = 2/9
:
12x^6 - 9x^4 + 6x^2 =
:
12(x^2)^3 - 9(x^2)^2 + 6(x^2) =
:
x^2(12(x^2)^2 -9(x^2) +6) =
:
2/9(12(2/9)^2 -9(2/9) +6) =
:
2/9((12*4/81) -2 +6) = 1.02
: